Colleges and universities in Country A offering graduate degrees enjoyed​ double-digit growth for nearly a​ decade, particularly from applicants of Country B. But applications from these students declined 5 percent for the last academic​ year, dropping the overall​ graduate-school application growth rate from foreign students to just 1 percent. This has schools that offer​ business, engineering, life​ sciences, and physical and earth sciences graduate degrees wondering how they will fill their classrooms. A bright spot is applications from Country​ C, which increased 24 percent because of a push by the its government to boost advanced degrees from Country A schools. Many schools in Country A are wondering if this market is worthy of more marketing efforts to replace the lost applicants from Country B. Develop a market potential estimation of this market.

Assume that there are approximately 27,500,000 citizens of Country C in the age​ 25-34. Only 15​% of the​ 25-34 year old age group already have graduate degrees and are not in the​ schools' target market. Among the remaining​ 25-34 year old age group 69​% can be considered as potential students because they completed at least upper secondary education. Also assume the average price for education in Country A is ​$180 thousand​ (for all years required to obtain a​ degree) and students will only earn one degree.

The number of buyers in the market is 16,128,750 ​(Round to the nearest whole​ number.)

The market demand is $__________ billion.
Answer #1

The number of potential customers in the country c = 27,500,000*(1-0.15)*.69 = 16,128,750

The market demand = potential cutomers*price for education = 16,128,750*180,000 = 2,903,175,000,000 = 2903.175 billion